### v1

@@ 3.2 Assistant Bot Functions
The assistant bot of AutoManager is responsible for users’ orders. Its main functions include helping users complete their orders, answering their questions, and recommending foods according to their preferences. The assistant bot’s logic is much more complex than an administrator bot’s because food ordering instructions at a fast food drive-through can be complex.
@@
LLM Parser and Ontology. Like the administrator bot, the assistant bot uses an LLM with instructions, predicate usages, and examples as prompts for knowledge extraction. In the assistant bot, the LLM extracts six predicates: need_recommend/2, order/2, specify/2, update/3, complete/0, and query/2. The predicate need_recommend(Content, Type) captures the user’s intention that asks for a recommendation. The ‘Type’ arity can be either ‘category’ or ‘upgrade’. Suppose the ‘Type’ is ‘category’. In that case, it means that the user needs the recommendation of a specific category of food, and then the ‘Content’ arity can be either ‘taco’, ‘burrito’, ‘quesadilla’, ‘nacho’, ‘bowl’, ‘vegan’, ‘chicken’, etc., or even ‘all’, indicating all categories of foods. If ‘Type’ is ‘upgrade’, indicating the user needs a recommendation for the popular toppings for the dish, then the ‘Content’ will be the name of the dish, such as ‘Crunchy Taco’, etc. The predicate order(Food, Number) captures the user’s intention of ordering, where the ‘Food’ arity is the name of a dish or combo that the user wants to order. The ‘Number’ arity indicates the quantity the user orders. The predicate specify(Combo, Dish) specifies a certain dish as the combo member, such as designating a specific taco for the taco bundle, or a drink for some combo. This predicate is used only for the combo. The predicate update(Dish, Operation, Option) captures the user’s intention of updating an ordered dish. The arity ‘Dish’ refers to the dish name, the ‘Operation’ arity refers to the upgrading operations, and the ‘Option’ refers to the corresponding options to choose from. In operations ‘change’, ‘add’, ‘no’, ‘less’, and ‘extra’, the option tells what ingredient to modify. For the operations ‘fresco’, ‘supreme’, and ‘grill’, the option can only be ‘yes’ or ‘no’, showing whether the user would like these upgrades. For the operation ‘size’, the option can be ‘regular’ or ‘large’. The predicate query(Category, Food) captures the user’s intention of asking. The ‘Category’ arity refers to the type of information the user asks for, while the ‘Food’ arity shows the object the user refers to. The ‘Food’ can also be ‘all’ in some cases. For example, the question “What are the available toppings?” can be expressed by the predicate query(‘add-on’, ‘all’), and the question “What’s the price of the Veggie Mexican Pizza?” can be expressed by the predicate query(‘price’, ‘Veggie Mexican Pizza’). When the user says that they have finished all they need to order, the predicate completed will be extracted.
@@
Ordering Strategy. After extracting predicates from the users’ input, the ASP-based goal-directed reasoning system decides what to say in the next round of dialogue. For the first several rounds of conversation, the assistant bot will listen to the customer’s requirements, and only reply by saying “I got it, what else do you want to order?” unless the user wants the bot to recommend food or answer questions. Once the user has indicated that he or she has completed all the dishes he or she wishes to order, where the predicate completed is extracted, the assistant bot, driven by CKTs, will check if any important information is missing and ask the customer corresponding questions. If the customer orders a single dish, the assistant bot will ask if they would like some toppings on the dish, and for the drink, the assistant will ask for the size, if not specified. If the order contains any combo, the CKT will go through each dish in the combo, asking the user to choose one optional dish from the listed choices, and then complete the CKT loop for this dish. Since the user may order the same dish repeatedly, the CKT will execute the same loop several times until all the dishes are complete. In this “ask” mode, the next-action predicate ask(Food, Option) or ask(Combo, Food, Option) is generated.
@@
Price Calculation. After collecting all the information, the assistant bot will calculate the price of the order. An extra charge will be applied if the order contains upgrades or add-ons. The output predicates, categorized as “check” mode, contain all the ordered dishes with their special requirements, grouped by each ordered dish, and the total price, as the example below shows.
@@
Question Answering and Recommendation. The question-answering function will be triggered if a query/2 predicate is extracted, and the assistant bot will look up an answer in the knowledge base. If the need_recommend/2 predicate is captured from the user input, the assistant bot will recommend dishes that satisfy the user’s requirements. Those mentioned in the previous conversation will not be recommended again. The “answer” mode produces the next-action predicate answer(Food, Property, Value) for question answering, and the “recommend” mode generates recommend(Type, Food) predicate to help the users find the food they would like.
@@
Response Generation. The response is generated using an LLM with some templates and prompts. In the “quit” mode, the only predicate output from the reasoning engine is quit/0, and the LLM will generate the response by rewriting the template for saying goodbye. Similarly, in the “else” mode, the LLM will respond by asking what else to help with. For the other modes, including “ask”, “recommend”, “answer”, and “check”, the response predicates contain a confirmation and the action part. The confirmation makes sure the LLM extracts the user’s intention correctly. For example, the extracted predicate update(‘Soft Taco’, ‘add’, ‘Cheese’). leads to the confirmation predicate confirm(‘add’, ‘Cheese’)., and generates the response like “Okay, you’d like to add more cheese.” in the LLM response generation step. The confirmation to the irrelevant predicate gives a response like “Sorry, I don’t understand. Let’s get back to your order.”, while the confirmation for unavailable dishes will be explained in Section 4. The output of the action part of each mode includes different predicates depending on the action mode, which is described in the above paragraphs. All the output predicates will be sent to an LLM for the natural language response generation. Like the knowledge extraction part, the prompt here also contains the meaning and usage of each predicate and a few examples.
4 Dual-Agent Collaboration
@@
Our AutoManager Chatbot system is designed as an administrator-assistant system managing a Taco Bell restaurant. In AutoManager, the administrator bot and the assistant bot collaborate on food sales and inventory management, where the administrator bot manages the inventory and the assistant bot interactively helps a customer place their order accordingly. As previously described, the dual system requires the two agents to collaborate on the tasks. Therefore, the dual system conversational agent should achieve the following functions: (i) Shareability: The two bots share the same knowledge base and can make changes to the knowledge base if necessary. In AutoManager, the administrator and assistant share the same menu, and the administrator can update the menu, such as adding new dishes or adjusting the price. The menu is part of the knowledge base. (ii) Independence: Without the other agent, Each bot itself should still run properly, even if the other bot is dormant, since each bot is designed for solving an independent task, and the only difference between the two bots is their target user. For example, in AutoManager, the administrator phase is only for restaurant employees, while the assistant interacts with restaurant customers. (iii) Collaboration: However, despite being independent, the two bots still constitute a dual system, meaning that they can communicate with each other through the knowledge base, which holds the key information for both bots to complete their respective tasks. In AutoManager, if the storage of lettuce, for instance, is low, the administrator will communicate this information to the assistant by updating the knowledge base, and the assistant will then reject the orders containing lettuce.
@@
Collaborative Rule Set (CRS): In AutoManager, the information about availability of ingredients (expressed by predicates runout/1 and restore/1) is captured from the dialogue by the administrator bot and saved in a temporary state file for the assistant bot to read. Apart from these predicates, some rules that reason on these predicates are also necessary for both bots: the order-taking bot would want to know what dishes are currently unavailable from the administrator bot and communicate to the customer. Both cases require the rules regarding the shortage of ingredients for the unavailable dishes to be explained, shared by the collaborative rule set. The CRS in AutoManager is listed below.
@@
Leveraging External Information For Internal Tasks (LEIFIT) Mechanism: With the help of CRS, the assistant bot can use the predicate unavailable/1 to complete the ordering task. This direct use of predicates in the CRS is called the LEIFIT mechanism. In Automanager, for example, this unavailable/1 predicate is used for rejecting user requirements that cannot be fulfilled and listing proper answers to the user’s questions and recommendation demands. For each round of conversation, the reasoning engine first checks if the dish or topping the user orders is available according to the CRS. If there is more than one unavailable order, the bot will generate a list of the unavailable dishes and the missing ingredient as the confirmation, e.g.,
@@
In the LEIFIT mechanism, we ensure each round of conversation updates the temporary information list once and checks the CRS consistency once for security reasons. The update on the knowledge base will only be executed after both bots end their conversations. The update operation is considered an atomic operation to avoid inconsistency during the update.
5 Implementation
@@
System Updating: In AutoManager, the administrator and assistant bot are required to maintain a list of the current state. However, the s(CASP) goal-directed ASP system does not support automatic state updates. Therefore, we constructed a general updating system for the bots.
@@
Then, the reasoning engine uses the predicates in both the state and the waiting list for the next action generation. Finally, all available predicates will be added to the state without the prefix. In the assistant bot, the case is more complicated. For example, if the customer wants to upgrade a dish he or she did not order previously, or the dish is no longer available, the assistant bot will reject adding it to the state by following the rules.
@@ 6 Conversation Samples
The assistant bot taking human customer order is shown next.
Assistant:
@@ 7 Evaluation
Call Option Administrator Assistant (1) Assistant (10)
@@
Table 1: Time cost for s(CASP) reasoning call and total response generation in the administrator bot and the assistant bot of AutoManager. Here “1” means the assistant bot is called with only one requirement predicate in the list, and “10” means 10 predicates are captured in the state list.
Efficiency Evaluation. We experimented on the processing time of the administrator bot and the assistant bot for each round of reply generation in Table 1. The average time consumption for 50 rounds is around 2.6 seconds for the administrator bot and 4.2 seconds for the assistant bot for 10 user requirements, which is an acceptable waiting time for serving sessions. The result also shows that the average time consumption for ASP reasoning is about 9.9 milliseconds for the administrator bot. For the assistant bot, the time consumption varies by the size of the user’s requirements. As the number of user requirements increased, it took longer for s(CASP) to reason. However, when the requirement number comes to 10, which is enough for an ordinary customer, the time spent on s(CASP) is only 886.5 milliseconds, which is a reasonable computation time. In addition, the time cost of the reasoning step is much shorter than the total time consumption, which also indicates that most of the time is spent calling GPT-4. We believe that LLMs with faster response times will emerge shortly, and the overall execution time will be further shortened.
@@
Table 2: Human Evaluation Results comparing our AutoManager with the real-world Taco Bell Drive-Thru AI system. We tested on five conversations with five human evaluators.
Human Evaluation. We also manually evaluated the conversation quality of our AutoManager and compared its results with the real-world Taco Bell drive-thru AI order-taking system. Note that the latter is a custom implementation watched over by a human. Since we do not have access to the Taco Bell drive-thru AI system, we collected the reported conversations through the Internet. We went through the videos posted online that include the entire conversation with Taco Bell AI and extracted transcripts from them. Then, we applied these conversations in our AutoManager system following the same requirements for the orders. As a result, we generated five pairs of conversations, where each pair contains one conversation from the Taco Bell drive-thru AI and one from AutoManager. We asked five well-educated graders to grade these five conversation pairs on seven metrics, following the human evaluation on the AutoConcierge by autoconcierge: Understanding, Truthfulness, Coherency, Fluency, Task Ease, Expected Behavior, and Total Satisfactory. The result can be found in Table 2. The result shows that AutoManager leads dominantly in understanding, task ease, and expected behavior, indicating that the core advantage of AutoManager over Taco Bell AI is its reliability and human-like behavior.
8 Conclusion and Future Work

### v2

@@ III-B Customer Service Bot Functions
The customer service bot of AutoManager is responsible for users’ orders. Its main functions include helping users complete their orders, answering their questions, and recommending foods according to their preferences. The customer service bot’s logic is much more complex than a manager bot’s because food ordering instructions at a fast food drive-through can be complex.
@@
LLM Parser and Ontology. Like the manager bot, the customer service bot uses an LLM with instructions, predicate usages, and examples as prompts for knowledge extraction. In the customer service bot, the LLM extracts six predicates: need_recommend/2, order/2, specify/2, update/3, complete/0, and query/2. The predicate need_recommend(Content, Type) captures the user’s intention that asks for a recommendation. The ‘Type’ arity can be either ‘category’ or ‘upgrade’. Suppose the ‘Type’ is ‘category’. In that case, it means that the user needs the recommendation of a specific category of food, and then the ‘Content’ arity can be either ‘taco’, ‘burrito’, ‘quesadilla’, ‘nacho’, ‘bowl’, ‘vegan’, ‘chicken’, etc., or even ‘all’, indicating all categories of foods. If ‘Type’ is ‘upgrade’, indicating the user needs a recommendation for the popular toppings for the dish, then the ‘Content’ will be the name of the dish, such as ‘Crunchy Taco’, etc. The predicate order(Food, Number) captures the user’s intention of ordering, where the ‘Food’ arity is the name of a dish or combo that the user wants to order. The ‘Number’ arity indicates the quantity the user orders. The predicate specify(Combo, Dish) specifies a certain dish as the combo member, such as designating a specific taco for the taco bundle, or a drink for some combo. This predicate is used only for the combo. The predicate update(Dish, Operation, Option) captures the user’s intention of updating an ordered dish. The arity ‘Dish’ refers to the dish name, the ‘Operation’ arity refers to the upgrading operations, and the ‘Option’ refers to the corresponding options to choose from. In operations ‘change’, ‘add’, ‘no’, ‘less’, and ‘extra’, the option tells what ingredient to modify. For the operations ‘fresco’, ‘supreme’, and ‘grill’, the option can only be ‘yes’ or ‘no’, showing whether the user would like these upgrades. For the operation ‘size’, the option can be ‘regular’ or ‘large’. The predicate query(Category, Food) captures the user’s intention of asking. The ‘Category’ arity refers to the type of information the user asks for, while the ‘Food’ arity shows the object the user refers to. The ‘Food’ can also be ‘all’ in some cases. For example, the question “What are the available toppings?” can be expressed by the predicate query(‘add-on’, ‘all’), and the question “What’s the price of the Veggie Mexican Pizza?” can be expressed by the predicate query(‘price’, ‘Veggie Mexican Pizza’). When the user says that they have finished all they need to order, the predicate completed will be extracted.
@@
Ordering Strategy. After extracting predicates from the users’ input, the ASP-based goal-directed reasoning system decides what to say in the next round of dialogue. For the first several rounds of conversation, the customer service bot will listen to the customer’s requirements, and only reply by saying “I got it, what else do you want to order?” unless the user wants the bot to recommend food or answer questions. Once the user has indicated that he or she has completed all the dishes he or she wishes to order, where the predicate completed is extracted, the customer service bot, driven by CKTs, will check if any important information is missing and ask the customer corresponding questions. If the customer orders a single dish, the service bot will ask if they would like some toppings on the dish, and for the drink, the service bot will ask for the size, if not specified. If the order contains any combo, the CKT will go through each dish in the combo, asking the user to choose one optional dish from the listed choices, and then complete the CKT loop for this dish. Since the user may order the same dish repeatedly, the CKT will execute the same loop several times until all the dishes are complete. In this “ask” mode, the next-action predicate ask(Food, Option) or ask(Combo, Food, Option) is generated.
@@
Price Calculation. After collecting all the information, the service bot will calculate the price of the order. An extra charge will be applied if the order contains upgrades or add-ons. The output predicates, categorized as “check” mode, contain all the ordered dishes with their special requirements, grouped by each ordered dish, and the total price, as the example below shows.
@@
Question Answering and Recommendation. The question-answering function will be triggered if a query/2 predicate is extracted, and the customer service bot will look up an answer in the knowledge base. If the need_recommend/2 predicate is captured from the user input, the service bot will recommend dishes that satisfy the user’s requirements. Those mentioned in the previous conversation will not be recommended again. The “answer” mode produces the next-action predicate answer(Food, Property, Value) for question answering, and the “recommend” mode generates recommend(Type, Food) predicate to help the users find the food they would like.
@@
Response Generation. The response is generated using an LLM with some templates and prompts. In the “quit” mode, the only predicate output from the reasoning engine is quit/0, and the LLM will generate the response by rewriting the template for saying goodbye. Similarly, in the “else” mode, the LLM will respond by asking what else to help with. For the other modes, including “ask”, “recommend”, “answer”, and “check”, the response predicates contain a confirmation and the action part. The confirmation makes sure the LLM extracts the user’s intention correctly. For example, the extracted predicate update(‘Soft Taco’, ‘add’, ‘Cheese’). leads to the confirmation predicate confirm(‘add’, ‘Cheese’)., and generates the response like “Okay, you’d like to add more cheese.” in the LLM response generation step. The confirmation to the irrelevant predicate gives a response like “Sorry, I don’t understand. Let’s get back to your order.”, while the confirmation for unavailable dishes will be explained in Section IV. The output of the action part of each mode includes different predicates depending on the action mode, which is described in the above paragraphs. All the output predicates will be sent to an LLM for the natural language response generation. Like the knowledge extraction part, the prompt here also contains the meaning and usage of each predicate and a few examples.
IV Dual-Agent Collaboration
@@
Our AutoManager Chatbot system is designed as a manager-customer-service system managing a Taco Bell restaurant. The manager bot and the customer service bot collaborate on food sales and inventory management, where the manager bot maintains the inventory and the service bot interactively helps customers place their orders accordingly. To maintain the system to work effectively, the two agents should achieve the following functions: (i) Shareability: The two bots share the same knowledge base and can make changes to it if necessary. In AutoManager, the manager and service bot share the same menu, and the manager bot can update it, such as adding new dishes or adjusting the price. The menu is part of the knowledge base. (ii) Independence: Without the other agent, each bot alone should still run properly, even if the other bot is dormant, since each bot is designed for solving an independent task, and the only difference between the two bots is their target user. For example, in AutoManager, the manager phase is only for restaurant employees, while the customer service phase interacts with restaurant customers. (iii) Collaboration: However, despite being independent, the two bots still constitute a dual system, meaning that they can communicate with each other through the knowledge base, which holds the key information for both bots to complete their respective tasks. In AutoManager, if the storage of lettuce, for instance, is low, the manager bot will communicate this information to the customer service bot by updating the knowledge base, and the service bot will then reject the orders containing lettuce.
@@
Shared Commonsense Knowledge and Rules: In AutoManager, the information about the availability of ingredients (expressed by predicates runout/1 and restore/1) is captured from the dialogue by the manager bot and saved in a temporary state file for the customer service bot to read. Apart from these predicates, some rules that reason on these predicates are also necessary for both bots: the order-taking bot would want to know what dishes are currently unavailable from the manager bot and communicate this to the customer. Both cases require the rules regarding the shortage of ingredients for the unavailable dishes to be explained, shared by the collaborative rule set. The shared rules in AutoManager are listed below.
@@
Leveraging External Information For Internal Tasks (LEIFIT) Mechanism: With commonsense knowledge and rules regarding unavailability, the service bot can use the predicate unavailable/1 to complete the ordering task. This direct use of shared predicates and rules for reasoning is called the LEIFIT mechanism. In Automanager, for example, this unavailable/1 predicate is used for rejecting user requirements that cannot be fulfilled and listing proper answers to the user’s questions and recommendation demands. For each round of conversation, the reasoning engine first checks if the dish or topping the user orders is available according to the shared rules. If there is more than one unavailable order, the bot will generate a list of the unavailable dishes and the missing ingredient as the confirmation, e.g.,
@@
In the LEIFIT mechanism, we ensure each round of conversation updates the temporary information list once and checks the consistency of the shared commonsense knowledge once for security reasons. The update on the knowledge base will only be executed after both bots end their conversations. The update operation is considered an atomic operation to avoid inconsistency during the update.
V Implementation
@@
System Updating: In AutoManager, the manager and service bot are required to maintain a list of the current state. However, the s(CASP) goal-directed ASP system does not support automatic state updates. Therefore, we constructed a general updating system for the bots.
@@
Then, the reasoning engine uses the predicates in both the state and the waiting list for the next action generation. Finally, all available predicates will be added to the state without the prefix. In the service bot, the case is more complicated. For example, if the customer wants to upgrade a dish he or she did not order previously, or the dish is no longer available, the assistant bot will reject adding it to the state by following the rules.
@@ VI Conversation Samples
Then the service bot took a human customer’s order.
ServiceBot:
@@ VII Evaluation
Call Option Manager Service (1) Service (10)
@@
TABLE I: Time cost for s(CASP) reasoning call and total response generation in the manager bot and the customer service bot of AutoManager. Here “1” means the service bot is called with only one requirement predicate in the list, and “10” means 10 predicates are captured in the state list.
Efficiency Evaluation. We experimented on the processing time of the manager and service bot for each round of reply generation in Table I. The average time consumption for 50 rounds is around 2.6 seconds for the administrator bot and 4.2 seconds for the assistant bot for 10 user requirements, which is an acceptable waiting time for serving sessions. The result also shows that the average time consumption for ASP reasoning is about 9.9 milliseconds for the administrator bot. For the assistant bot, the time consumption varies by the size of the user’s requirements. As the number of user requirements increased, it took longer for s(CASP) to reason. However, when the requirement number comes to 10, which is enough for an ordinary customer, the time spent on s(CASP) is only 886.5 milliseconds, which is a reasonable computation time. In addition, the time cost of the reasoning step is much shorter than the total time consumption, which also indicates that most of the time is spent calling GPT-4. We believe that LLMs with faster response times will emerge shortly, and the overall execution time will be further shortened.
@@
TABLE II: Human Evaluation Results comparing our AutoManager with the real-world Taco Bell Drive-Thru AI system. We tested on five conversations with five human evaluators.
Human Evaluation. We also manually evaluated the conversation quality of our AutoManager and compared its results with the real-world Taco Bell drive-thru AI order-taking system. Note that the latter is a custom implementation watched over by a human. Since we do not have access to the Taco Bell drive-thru AI system, we collected the reported conversations through the Internet. We went through the videos posted online that include the entire conversation with Taco Bell AI and extracted transcripts from them. Then, we applied these conversations in our AutoManager system following the same requirements for the orders. As a result, we generated five pairs of conversations, where each pair contains one conversation from the Taco Bell drive-thru AI and one from AutoManager. We asked five well-educated graders to grade these five conversation pairs on seven metrics, following the human evaluation on the AutoConcierge by [3]: Understanding, Truthfulness, Coherency, Fluency, Task Ease, Expected Behavior, and Total Satisfactory. The result can be found in Table II. The result shows that AutoManager leads dominantly in understanding, task ease, and expected behavior, indicating that the core advantage of AutoManager over Taco Bell AI is its reliability and human-like behavior.
VIII Conclusion and Future Work